### v2

# Rankify: A Comprehensive Python Toolkit for Retrieval, Re-Ranking, and Retrieval-Augmented Generation

Abstract.
Retrieval, re-ranking, and retrieval-augmented generation (RAG) are critical components of modern natural language processing (NLP) applications in information retrieval, question answering, and knowledge-based text generation. However, existing solutions are often fragmented, lacking a unified framework that easily integrates these essential processes. The absence of a standardized implementation, coupled with the complexity of retrieval and re-ranking workflows, makes it challenging for researchers to compare and evaluate different approaches in a consistent environment. While existing toolkits such as Rerankers and RankLLM provide general-purpose reranking pipelines, they often lack the flexibility required for fine-grained experimentation and benchmarking. In response to these challenges, we introduce Rankify, a powerful and modular open-source toolkit designed to unify retrieval, re-ranking, and RAG within a cohesive framework. Rankify supports a wide range of retrieval techniques, including dense and sparse retrievers, while incorporating state-of-the-art re-ranking models to enhance retrieval quality. Additionally, Rankify includes a collection of pre-retrieved datasets to facilitate benchmarking, available at Huggingface [^1] https://huggingface.co/datasets/abdoelsayed/reranking-datasets-light. To encourage adoption and ease of integration, we provide comprehensive documentation [^2] http://rankify.readthedocs.io/, an open-source implementation on GitHub [^3] https://github.com/DataScienceUIBK/rankify, and a PyPI package for effortless installation [^4] https://pypi.org/project/rankify/. By providing a unified and lightweight framework, Rankify allows researchers and practitioners to advance retrieval and re-ranking methodologies while ensuring consistency, scalability, and ease of use.
Neural IR, Dense Retrieval, Sparse Retrieval, Re-ranking, Toolkit
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Evaluation of retrieval resultsccs: Information systems Presentation of retrieval resultsccs: Information systems Top-k retrieval in databasesccs: Information systems Rank aggregation
Figure 1. Rankify logo.
1. Introduction
Information retrieval (IR) systems (Singhal et al., 2001; Chowdhury, 2010) are fundamental to many natural language processing (NLP) applications (Wang et al., 2024; Liu et al., 2024), including question-answering (Khamnuansin et al., 2024; Abdallah et al., 2024b), search engines (Croft et al., 2010; Xiong et al., 2024), and knowledge-based generation (Long et al., 2024). These systems often rely on a two-stage pipeline: a retriever that efficiently identifies a set of candidate documents and a re-ranker that refines these results to maximize relevance to the query (Yates et al., 2021; Abdallah et al., 2024a; Nogueira et al., 2019; Ren et al., 2021). This approach has proven highly effective, with retrieval and re-ranking methods achieving state-of-the-art performance across diverse NLP benchmarks.
Retrievers form the backbone of information retrieval systems by identifying a subset of documents relevant to a user’s query. The retrieval landscape includes sparse, dense, and hybrid methods. Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching by representing queries and documents as high-dimensional sparse vectors. These methods are highly effective when query terms closely align with document content but struggle to capture semantic relationships. Dense retrievers, including models like DPR (Yates et al., 2021), BPR (Yamada et al., 2021), ColBERT (Santhanam et al., 2021), BGE (Chen et al., 2023), Contriever (Izacard et al., 2021) and ANCE (Xiong et al., 2021a), overcome this limitation by encoding queries and documents into low-dimensional dense vectors using neural networks, enabling retrieval based on semantic similarity even when lexical overlap is minimal. However, dense retrieval requires significant computational resources for training and inference. Hybrid retrieval methods (Yang et al., 2019; Chen et al., 2022) integrate the strengths of both approaches, combining the precision of sparse retrievers with the semantic understanding of dense models to achieve a more balanced and robust retrieval system.
Toolkit # Datasets # Retrievers # Re-rankers # RAG Methods Modular Design Automatic Evaluation Corpus
FlashRAG (Jin et al., 2024) 32 3 2 12
FastRAG (Abane et al., 2024) 0 1 3 7
AutoRAG (Kim et al., 2024a) 4 2 7 1
LocalRQA (Yu et al., 2024) 0 2 0 0
Rerankers (Clavié, 2024) 0 0 15 0
Rankify 40 7 24 3+
Table 1. Comparison with other Retriever, Re-ranking, and RAG toolkits. Modular Component refers to whether the toolkit is composed of modular components. Automatic Evaluation indicates if the toolkit provides automated evaluation capabilities to assess the performance of various datasets. Corpus Helper shows if the toolkit offers auxiliary tools for processing corpus, including cleaning and chunking.
Re-ranking techniques enhance the initial retrieval results by ensuring the most relevant documents appear at the top. These methods are categorized into three main approaches: (1) Pointwise reranking (Nogueira et al., 2019; Abdallah et al., 2025b, a; Sachan et al., 2022) treats reranking as a regression or classification task, assigning independent relevance scores to each document. (2) Pairwise reranking (Huang and Chen, 2024; Qin et al., 2023; Sinhababu et al., 2024) refines ranking by comparing document pairs and optimizing their relative order based on relevance. However, this approach treats all document pairs equally, sometimes improving lower-ranked results at the expense of top-ranked ones. (3) Listwise reranking (Sun et al., 2023a; Pradeep et al., 2023b; Yoon et al., 2024) considers the entire document list, prompting LLMs with the query and a subset of candidates for reranking. Due to input length limitations, these methods tend to employ a sliding window strategy, progressively refining rankings from back to front.
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Mozafari et al., 2024; Abdallah and Jatowt, 2023; Gao et al., 2023) enhances language models by integrating retrieval and generation, making them more effective in knowledge-intensive tasks. Rather than relying solely on pre-trained knowledge, RAG dynamically retrieves relevant documents from external sources during inference, incorporating them into the generation process.
The growing complexity and diversity of retrieval, re-ranking, and RAG methods pose significant challenges for benchmarking, reproducibility, and integration. Existing toolkits, such as Pyserini (Lin et al., 2021), Rerankers (Clavié, 2024) and RankLLM (Pradeep et al., 2023a), often lack flexibility, enforce rigid implementations, and require extensive preprocessing, making them less suitable for research-driven experimentation. Additionally, retrieval and re-ranking datasets are scattered across different sources, complicating evaluation and comparison. To address these challenges, we introduce Rankify, an open-source framework that unifies retrieval, re-ranking, and RAG into a modular and extensible ecosystem with its logo shown in Figure 1. Rankify supports diverse retrieval techniques, integrates state-of-the-art re-ranking models, and provides curated pre-retrieved datasets to streamline experimentation. Designed with flexibility, it enables researchers to efficiently build, evaluate, and extend retrieval pipelines while ensuring consistency in benchmarking.
The main contributions of this paper are as follows:
  • Curated retrieval datasets and precomputed embeddings: Rankify provides 40 datasets with 1,000 pre-retrieved documents per query across various domains (QA, dialogue, entity linking, etc.). It also includes pre-computed Wikipedia and MS MARCO corpora for multiple retrievers, eliminating preprocessing overhead.
  • Diverse retriever and re-ranking support: Rankify integrates dense (DPR, ANCE, BPR, ColBERT, BGE, Contriever) and sparse (BM25) retrievers, along with 24 state-of-the-art re-ranking models, enabling seamless retrieval and ranking experimentation.
  • RAG integration and evaluation tools: Rankify bridges retrieval, re-ranking, and RAG by passing retrieved documents to LLMs for evaluation.
  • Comprehensive evaluation tools: Rankify offers a diverse range of evaluation metrics and tools for retrieval and question answering. The framework is accompanied by extensive online documentation, making it easy for users to explore its features. Additionally, Rankify is freely available on PyPI and GitHub, ensuring accessibility for researchers and practitioners.
2. Related Work
2.1. Retrievers
Retrieval models have evolved significantly, transitioning from traditional sparse retrievers to sophisticated dense and hybrid retrieval techniques. Sparse retrievers, such as BM25 (Robertson et al., 1995), remain widely used due to their strong lexical matching capabilities and generalization. However, they struggle with capturing semantic relationships, leading to the rise of dense retrieval methods (Lee et al., 2019), which leverage pre-trained neural encoders to generate embeddings for queries and documents. Notable advancements, such as DPR (Yates et al., 2021) and ANCE (Xiong et al., 2021a), improve retrieval effectiveness by employing contrastive training with in-batch negatives (Gillick et al., 2019) and hard-negative mining (Karpukhin et al., 2020; Xiong et al., 2021b). Multi-vector approaches like ColBERT (Khattab and Zaharia, 2020) enable fine-grained interaction between queries and documents, albeit at the cost of increased computational overhead. To address the limitations of dense retrievers, hybrid models (Gao et al., 2021; Ma et al., 2022) combine sparse and dense approaches, utilizing both term matching and semantic understanding for enhanced retrieval performance.
Recent research has also explored retrieval enhancements through knowledge distillation, curriculum learning, and pre-training. Distillation techniques transfer knowledge from cross-encoders or teacher retrievers to student models, improving dense retriever effectiveness (Qu et al., 2021; Chen et al., 2021). Curriculum learning strategies, which progressively train retrievers from easy to hard samples (Zeng et al., 2022; Lin et al., 2022a), have been shown to enhance both supervised and zero-shot retrieval effectiveness. Meanwhile, pre-training methods such as contrastive learning (Izacard et al., 2021) and masked autoencoding (Gao and Callan, 2021; Liu and Shao, 2022) have significantly improved the generalization of retrieval models. Additionally, approaches like query augmentation (Mao et al., 2021; Ma et al., 2021) generate synthetic training data to enhance retrieval robustness.
2.2. Re-Ranking Methods
Re-ranking plays a crucial role in refining retrieved document lists to improve ranking quality. Traditional approaches include pointwise learning-to-rank (LTR) methods, where each document is scored independently (Zhu et al., 2021; Cinar and Renders, 2020), and pairwise methods, which optimize the relative ordering of document pairs (Cao et al., 2007; Jia et al., 2021). Listwise approaches, such as LambdaRank and ListNet, directly optimize ranking metrics like Discounted Cumulative Gain (DCG) to enhance top-ranked results (Burges, 2010; Liu et al., 2017; Cao et al., 2007). More recently, deep neural networks and large language models (LLMs) have revolutionized re-ranking, enabling context-aware ranking adjustments. Cross-encoder models, which jointly encode queries and candidate documents, have demonstrated significant improvements in ranking accuracy by capturing fine-grained interactions between them. Additionally, zero-shot re-ranking approaches leverage models like GPT-4 for listwise re-ranking without task-specific training (Achiam et al., 2023; Sun et al., 2023c), while knowledge distillation techniques such as RankZephyr fine-tune smaller models to approximate LLM performance with reduced computational cost (Pradeep et al., 2023b).
Re-ranking is also increasingly integrated into retrieval-augmented generation (RAG) systems, where refining retrieved documents before response generation ensures higher relevance and coherence in generated outputs. These advancements highlight the transition from traditional ranking techniques to sophisticated, model-driven re-ranking approaches that enhance retrieval quality in diverse NLP applications. Recent innovations in re-ranking include leveraging large language models (LLMs) such as GPT-4 (Achiam et al., 2023) for zero-shot and listwise re-ranking (Sun et al., 2023c). Other approaches, such as RankZephyr (Pradeep et al., 2023b), utilize knowledge distillation to fine-tune smaller models based on LLM outputs, enabling efficient and effective re-ranking.
2.3. Retrieval-Augmented Generation (RAG)
Retrieval-Augmented Generation (RAG) combines retrieval and generation to enhance language models in knowledge-intensive tasks. Early RAG systems (Lewis et al., 2020) retrieve context documents relevant to a query and use these as input for generation, enabling models to answer queries beyond their training data. More recent approaches explore the integration of retrieval with both encoder-decoder (Izacard and Grave, 2020) and decoder-only (Raffel et al., 2020) architectures, optimizing performance for a variety of downstream tasks.
One key area of focus in RAG is mitigating hallucinations and improving the factuality of generated responses (Manakul et al., 2023). Techniques such as self-consistency (Wang et al., 2022b) and noisy context filtering (Fang et al., 2024) have been proposed to enhance the reliability of RAG outputs. However, the quality of retrieved documents remains a critical factor in the overall effectiveness of RAG pipelines. Recent works emphasize the importance of combining retrieval with robust re-ranking techniques to ensure that the retrieved context is both relevant and precise (Kamalloo et al., 2023).
2.4. Frameworks
Several frameworks have been developed to support research and development in retrieval and re-ranking systems. rerankers (Clavié, 2024) is a lightweight Python library that provides a unified interface to commonly used re-ranking approaches, simplifying the integration of various methods into retrieval pipelines. Similarly, RankLLM [^5] https://github.com/castorini/rank_llm/tree/main offers a Python toolkit focused on reproducible information retrieval research, emphasizing listwise re-ranking with large language models. Additionally, FlashRAG (Jin et al., 2024) and AutoRAG (Kim et al., 2024b) introduce a modular toolkit designed for efficient retrieval-augmented generation (RAG) research, aiming to enhance flexibility and extensibility in experimental settings. These frameworks, along with others like LangChain (Chase, 2022), LlamaIndex (Liu, 2022), and DSPy (Khattab et al., 2023), provide researchers and practitioners with versatile tools to develop and evaluate retrieval and RAG applications. A comparison of retrieval, re-ranking, and RAG toolkits is presented in Table 1, highlighting the advantages of Rankify in dataset diversity, retriever and re-ranker support, and modularity.
Figure 2. The architecture of the Rankify, showing the interplay between its core modules: Datasets, Retrievers, Re-Rankers, and RAG Evaluation. Each module operates independently while seamlessly integrating with others, enabling end-to-end retrieval and ranking workflows.
3. Rankify
Modern information retrieval systems rely on a combination of retrieval, ranking, and generative techniques to surface relevant content and generate accurate responses. However, existing solutions are often fragmented, requiring researchers and practitioners to integrate multiple toolkits to experiment with different retrieval and ranking strategies. Rankify addresses this challenge by offering an end-to-end retrieval and ranking ecosystem, enabling seamless experimentation across a wide range of retrieval, re-ranking, and retrieval-augmented generation (RAG) models.
Unlike existing retrieval and ranking frameworks that focus on a single aspect of the pipeline, Rankify is designed to be modular, extensible, and lightweight, allowing users to plug in different retrievers, ranking models, and RAG techniques with minimal effort. Built in Python, it leverages machine learning libraries such as TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2019), Spacy (Honnibal et al., 2020), and Pyserini (Lin et al., 2021) . The framework is available on PyPI [^6] https://pypi.org/project/rankify/, making installation straightforward:
\$ pip install rankify
To ensure accessibility and reproducibility, Rankify supports both precomputed and real-time retrieval workflows. Users can work with pre-retrieved datasets for benchmarking or perform live document retrieval and re-ranking on large-scale corpora. The framework includes ready-to-use indices for Wikipedia and MS MARCO, eliminating the need for time-consuming indexing. Additionally, it supports different retrievers, giving users the flexibility to test different retrieval paradigms without modifying their pipeline.
Rankify consists of four core modules: Datasets, Retrievers, Re-Rankers, and RAG. The Datasets module provides standardized access to benchmark datasets, including Natural Questions, TriviaQA, and HotpotQA, as well as the ability to create custom datasets. The Retrievers module integrates state-of-the-art retrieval methods, including BM25, DPR, ANCE, BGE, and ColBERT, allowing users to experiment with different retrieval strategies. The Re-Rankers module includes 24 ranking models with 41 sub-methods, supporting pointwise, pairwise, and listwise re-ranking techniques. Finally, the RAG module enables retrieval-augmented generation using models such as LLAMA, GPT, and T5, supporting in-context learning, Fusion-in-Decoder (FiD), and retrieval-augmented latent memory (RALM).
A key advantage of Rankify is its unified API, which abstracts away implementation details, allowing users to experiment with different models and configurations using a simple interface. Figure 2 illustrates the architecture of Rankify, demonstrating how retrieval, ranking, and generation components interact within the pipeline. The framework is designed to be scalable and adaptable, making it suitable for both academic research and real-world applications in search, question answering, and document ranking.
To encourage collaboration and community contributions, Rankify is open-source and actively maintained [^7] https://github.com/abdoelsayed2016/reranking. Extensive documentation [^8] http://rankify.readthedocs.io/ and inline docstrings provide clear guidance on using the framework, ensuring ease of adoption for both beginners and experienced researchers. The next sections explore its capabilities in greater detail, covering retrieval benchmarks, ranking evaluations, and retrieval-augmented generation results.
Figure 3. Summary of datasets used in our Rankify. The figure illustrates the dataset sizes across different categories. The bubble size represents the total number of samples (Train + Dev + Test) for each dataset. Categories include QA, Multi-Hop QA, Long-Form QA, Multiple-Choice, Entity-Linking, Slot Filling, Fact Verification, Dialog Generation, Summarization, and Other specialized datasets. The log scale is used for better visualization of dataset size variations.
Figure 4. Schema of the Dataset class, illustrating the objects used to represent a dataset in Rankify.
3.1. Prebuilt Retrieval Corpora and Indexes
Rankify provides two large-scale retrieval corpora: Wikipedia (Karpukhin et al., 2020) and MS MARCO (Nguyen et al., 2017), enabling researchers to conduct retrieval experiments on well-established benchmarks. To support efficient retrieval, Rankify includes pre-builtt index files for each retriever, allowing users to directly query the corpora without requiring costly indexing. Each corpus has been indexed for multiple retrieval models, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT. By providing precomputed indexes, Rankify simplifies large-scale retrieval research while maintaining consistency across different retrieval pipelines. Users can view all available datasets and retrievers using the following command (see Listing 1):
Listing 1: This script retrieves the latest information about the available datasets, and displays metadata for each dataset in the terminal.
1from rankify.dataset.dataset import Dataset
2Dataset.avaiable_dataset()
3.2. Datasets
With Rankify, we aim to provide a standardized and efficient way to handle datasets for retrieval, re-ranking, and retrieval-augmented generation (RAG). Given the complexity of retrieval workflows, Rankify integrates pre-retrieved datasets with structured annotations, allowing users to experiment with different retrieval and re-ranking methods using just a few lines of code. Our dataset design unifies multiple annotation styles and storage formats, enabling a seamless interface for loading, processing, and benchmarking retrieval-based tasks.
The datasets included in Rankify fall into multiple categories, as illustrated in Figure 3, which showcases their distribution based on task type and dataset size. Each dataset in Rankify follows a standardized schema that consists of three main components: (1) Question: The input query for which relevant information needs to be retrieved. (2) Answers: The expected correct responses associated with the question. (3) Contexts (Or Retrieved Documents): A ranked list of candidate documents retrieved by a specific retriever. The Dataset class in Rankify encapsulates these attributes and provides easy access to retrieval, evaluation, and benchmarking. The dataset schema is depicted in Figure 4, illustrating how questions, answers, and contexts interact within Rankify.
To facilitate benchmarking and ease of use, Rankify provides pre-retrieved datasets where each query is associated with 1,000 retrieved documents from different retrievers. These datasets allow users to test different re-ranking and retrieval models without the need for time-consuming indexing.
As mentioned above, to ensure standardized retrieval benchmarking, Rankify provides pre-retrieved datasets, where each query is associated with 1,000 retrieved documents from Wikipedia. These datasets span multiple retrieval tasks, including question answering, multi-hop reasoning, temporal retrieval, entity linking, fact verification, dialogue generation, and summarization. To enable fair comparisons across different retrieval methods, each dataset has been processed using multiple retrievers, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT. By applying the same retrieval pipeline across all the datasets, Rankify eliminates the need for computationally expensive indexing while maintaining consistency in evaluation.
Rankify allows users to load and explore datasets with minimal effort. Below is an example demonstrating how to download and inspect a dataset using Rankify ’s API:
Listing 2: Loading a dataset in Rankify and inspecting its structure.
1from rankify.dataset.dataset import Dataset
2# Load the nq-test dataset retrieved with BM25
3dataset = Dataset(retriever="bm25", dataset_name="nq-test")
4documents = dataset.download(force_download=False)
5# Inspect the first document
6print(documents[0])
Beyond preloaded datasets, Rankify enables users to define their own datasets using the Dataset class. Below is an example demonstrating how to create a custom retrieval dataset:
Listing 3: Creating a custom dataset in Rankify.
1from rankify.dataset.dataset import Dataset, Document, Question, Answer, Context
2# Define a new dataset with a single question and answer
3question = Question("What is the capital of France?")
4answer = Answer(["Paris"])
5context = [Context(score=0.9, has_answer=True, id=1, title="France", text="The capital is Paris.")]
6# Create the document and dataset
7document = Document(question, answer, context)
8custom_dataset = Dataset(retriever="custom", dataset_name="MyDataset")
9custom_dataset.documents = [document]
10# Save the dataset
11custom_dataset.save_dataset("my_dataset.json")
Users can load their own datasets using load_dataset and load_dataset_qa functions. The load_dataset function is designed for structured datasets that include both queries and retrieved documents. On the other hand, load_dataset_qa is intended for question-answering datasets that contain only queries and answers, allowing users to use them directly for retrieval tasks. The following examples demonstrate how users can load these datasets in Rankify:
Listing 4: Loading a dataset with documents and a QA-only dataset.
1from rankify.dataset.dataset import Dataset
2# Load a dataset that includes retrieved documents
3retrieval_dataset = Dataset.load_dataset("path/to/retrieval_dataset.json")
4# Load a QA dataset with only questions and answers
5qa_dataset = Dataset.load_dataset_qa("path/to/qa_dataset.jsonl")
3.3. Retriever Models
Rankify supports a diverse set of retriever models to facilitate document retrieval for various information retrieval tasks. The current version includes both sparse and dense retrievers. Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching, while dense retrievers, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021a), and BGE (Chen et al., 2023), and ColBERT (Khattab and Zaharia, 2020).
The retrievers implemented in Rankify are accessible via a unified interface, where users can specify the retrieval method and parameters. The available retriever models include BM25, DPR, ANCE, Contriever (Izacard et al., 2021), BGB, and ColBERT. Rankify allows users to retrieve documents efficiently using different retrieval models. The retrievers can be applied to any dataset supported by Rankify, making it flexible for experimentation and benchmarking. Users can easily initialize a retriever and apply it to a list of query-document pairs, as demonstrated in Listing 5.
Listing 5: Using a retriever model in Rankify to retrieve documents.
1from rankify.dataset.dataset import Document, Question, Answer
2from rankify.retrievers.retriever import Retriever
3
4# Define sample documents
5documents = [
6 Document(question=Question("the cast of a good day to die hard?"),
7 answers=Answer(["Jai Courtney", "Bruce Willis"]), contexts=[]),
8 Document(question=Question("Who wrote Hamlet?"),
9 answers=Answer(["Shakespeare"]), contexts=[])
10]
11# Initialize a retriever (example: ColBERT)
12retriever = Retriever(method="colbert", model="colbert-ir/colbertv2.0", n_docs=5, index_type="msmarco")
13# Retrieve documents
14retrieved_documents = retriever.retrieve(documents)
15# Print retrieved documents
16for i, doc in enumerate(retrieved_documents):
[…]
18 print(doc)
In Listing 5, users can specify the retrieval corpus, select the desired retriever model, and define the number of documents to retrieve. This flexibility allows researchers to experiment with different retrieval settings, optimizing their retrieval strategy based on the task requirements.
3.4. Re-Ranking Models
In information retrieval, two-stage pipelines are widely used to maximize retrieval performance. The first stage involves retrieving a set of candidate documents using a computationally efficient retriever, followed by a second stage where these documents are re-ranked using a stronger, typically neural network-based, model. This re-ranking step enhances retrieval quality by considering deeper semantic relationships between the query and documents. While effective, re-ranking models vary significantly in their architectures, trade-offs, and implementations, making it challenging for users to select the most suitable method for their specific needs.
To address this, Rankify provides a unified interface for re-ranking, enabling users to effortlessly switch between different models with minimal modifications. The framework supports a diverse set of 24 primary re-ranking models with 41 sub-methods, spanning different re-ranking strategies such as pointwise, pairwise, and listwise approaches. The implemented models include MonoBERT (Nogueira and Cho, 2019), MonoT5 (Nogueira et al., 2020), RankT5 (Zhuang et al., 2023), ListT5 (Yoon et al., 2024), ColBERT (Santhanam et al., 2021), RankGPT (Sun et al., 2023b), and various transformer-based re-rankers.
Users can apply these models to re-rank retrieved documents using a simple interface, as demonstrated in Listing 6.
Listing 6: Applying re-ranking in Rankify.
1from rankify.dataset.dataset import Document, Question, Answer
2from rankify.retrievers.retriever import Retriever
3from rankify.rerankers.reranker import Reranker
4# Define sample documents
5documents = [
6 Document(question=Question("the cast of a good day to die hard?"),
7 answers=Answer(["Jai Courtney", "Bruce Willis"]), contexts=[]),
8 Document(question=Question("Who wrote Hamlet?"),
9 answers=Answer(["Shakespeare"]), contexts=[])
10]
11# Initialize a retriever
12retriever = Retriever(method="colbert", model="colbert-ir/colbertv2.0", n_docs=5, index_type="msmarco")
13retrieved_documents = retriever.retrieve(documents)
14# Initialize a re-ranker
15reranker = Reranker(method="monot5",model_name="monot5-base")
16# Apply re-ranking
17reranked_documents = reranker.rerank(retrieved_documents)
18# Print re-ranked documents
19for i, doc in enumerate(reranked_documents):
[…]
21 print(doc)
Rankify supports both pre-trained models from Hugging Face and API-based re-ranking models such as RankGPT, Cohere, and Jina Reranker. Users can select from a variety of models, including cross-encoders, T5-based ranking models, and late-interaction retrieval models like ColBERT. The flexibility to switch between different re-ranking strategies allows researchers and practitioners to optimize ranking performance based on their task requirements.
3.5. Retrieval-Augmented Generation (RAG) Models
Retrieval-Augmented Generation (RAG) enhances language models by integrating retrieval mechanisms, allowing them to generate responses based on dynamically retrieved documents rather than relying solely on pre-trained knowledge. This approach is particularly effective for knowledge-intensive tasks such as open-domain question answering, fact verification, and knowledge-based text generation. Rankify provides a modular and extensible interface for applying multiple RAG methods, including zero-shot generation, Fusion-in-Decoder (FiD) (Izacard and Grave, 2020), and in-context learning  (Ram et al., 2023).
In Rankify, the Generator module enables seamless integration of RAG techniques, allowing users to experiment with different generative approaches. Users can specify the desired RAG method and model, applying generation strategies across retrieved documents.
Users can apply these methods to generate responses based on retrieved documents. Listing 7 demonstrates how to use Rankify’s RAG module with an in-context learning approach.
Listing 7: Applying Retrieval-Augmented Generation (RAG) in Rankify.
1from rankify.dataset.dataset import Document, Question, Answer, Context
2from rankify.generator.generator import Generator
3# Sample question and contexts
4question = Question("What is the capital of France?")
5answers = Answer(["Paris"])
6contexts = [
7 Context(id=1, title="France", text="The capital of France is Paris.", score=0.9),
8 Context(id=2, title="Germany", text="Berlin is the capital of Germany.", score=0.5)
9]
10# Create a Document
11doc = Document(question=question, answers=answers, contexts=contexts)
12# Initialize Generator with In-Context RALM
13generator = Generator(method="in-context-ralm", model_name=’meta-llama/Llama-3.1-8B’)
14# Generate answer
15generated_answers = generator.generate([doc, doc])
16print(generated_answers)
Rankify allows users to leverage large-scale language models such as LLaMA (Touvron et al., 2023), GPT-4 (Achiam et al., 2023), and T5-based models (Raffel et al., 2020) for retrieval-augmented generation. By supporting both encoder-decoder architectures (FiD (Izacard and Grave, 2020)) and decoder-only models (e.g., GPT, LLaMA), Rankify provides flexibility for optimizing generation quality based on task-specific requirements.
3.6. Evaluation Metrics
Rankify provides evaluation metrics for retrieval, re-ranking, and retrieval-augmented generation (RAG). For retrieval and re-ranking, Top-k accuracy measures whether relevant documents appear within the top-k results, allowing performance comparison before and after re-ranking. Listing 8 demonstrates how to compute Top-k accuracy.
Listing 8: Computing Top-k accuracy for retrieval and re-ranking in Rankify.
1from rankify.metrics.metrics import Metrics
2metrics = Metrics(documents)
3# Compute Top-k accuracy before and after re-ranking
4before_ranking = metrics.calculate_retrieval_metrics(ks=[1,5,10,20,50,100], use_reordered=False)
5after_ranking = metrics.calculate_retrieval_metrics(ks=[1,5,10,20,50,100], use_reordered=True)
6print("Before Ranking:", before_ranking)
7print("After Ranking:", after_ranking)
For QA and RAG, Rankify supports Exact Match (EM), recall, precision, and containment to evaluate generated answers. Listing 9 shows how these metrics are computed.
Listing 9: Computing QA and RAG evaluation metrics in Rankify.
1from rankify.metrics.metrics import Metrics
2qa_metrics = Metrics(documents)
3qa_results = qa_metrics.calculate_generation_metrics(generated_answers)
4print(qa_results)
By providing a unified evaluation module, Rankify ensures consistent benchmarking across retrieval, re-ranking, and RAG tasks.
4. Experimental Result and Discussion
4.1. Experiment Setup
Rankify enables researchers to benchmark retrieval, re-ranking, and RAG methods, evaluate their own approaches, and explore optimizations within these tasks. To demonstrate its capabilities, we conducted multiple experiments to provide reproducible benchmarks and performance insights. All main experiments were conducted on 8 NVIDIA A100 GPUs.
Experiments were performed on a diverse set of datasets, covering open-domain and multi-hop QA, fact verification, and temporal retrieval tasks. We evaluated retrieval performance on datasets such as Natural Questions (NQ), TriviaQA, HotpotQA, 2WikiMultiHopQA, and ArchivalQA, while re-ranking performance was analyzed on MSMARCO, WebQuestions, and PopQA. For retrieval evaluation, we measured Top-k accuracy at k=1, 5, 10, 20, 50, 100. Exact match (EM), Precision, Recall, Contains, and F1 were used as the primary evaluation metric for QA tasks.
We conducted experiments across all supported retrieval, re-ranking, and RAG methods. Retrieval models BM25, DPR, ANCE, BGE, and ColBERT. Re-ranking methods tested included MonoT5, RankT5, RankGPT, and various transformer-based re-rankers. For RAG, we evaluated zero-shot generation, Fusion-in-Decoder (FiD), in-context learning (RALM). These experiments highlight the adaptability of Rankify for benchmarking retrieval, ranking, and knowledge-grounded text generation.
Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%) Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%)
2WikiMultiHopQA (Ho et al., 2020) dev 17.17 31.72 40.14 47.73 55.88 61.17 ArchivalQA (Wang et al., 2022a) dev 18.18 33.29 39.72 45.96 53.68 58.74
train 8.85 19.78 28.15 38.16 51.16 57.70 test 17.68 32.55 39.26 45.36 52.76 58.08
ChroniclingAmericaQA (Piryani et al., 2024) test 4.18 11.23 15.83 20.87 27.49 32.48 AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) train 27.30 50.67 59.63 67.63 75.68 79.87
dev 4.26 11.80 16.42 21.50 28.19 33.15 dev 30.22 54.25 65.58 72.98 80.47 84.92
ARC (Clark et al., 2018) test 4.96 19.84 33.57 51.94 79.09 91.80 TriviaQA (Joshi et al., 2017) dev 48.70 67.57 72.84 77.38 81.19 83.59
dev 4.83 19.45 30.84 52.13 79.06 91.83 test 48.22 67.44 72.79 77.28 81.42 83.87
train 4.45 18.16 31.39 51.39 78.19 91.36 train 47.87 67.22 72.90 77.15 81.30 83.64
BoolQ (Clark et al., 2019) train 1.71 4.33 7.23 12.03 24.56 37.76 CommenseQA (Talmor et al., 2019) train 2.85 8.74 13.36 19.28 28.87 36.54
dev 1.22 3.94 6.61 11.68 23.67 37.49 dev 2.13 6.96 11.88 17.77 27.11 33.82
CuratedTREC (Lee et al., 2020) train 16.98 29.77 33.95 36.74 40.00 40.93 AY2 (Hoffart et al., 2011) train 23.81 41.60 48.84 55.88 63.78 69.62
test 18.14 31.63 34.65 38.84 41.86 43.02 dev 25.02 41.14 48.10 55.39 64.19 70.55
FEVER (Thorne et al., 2018; Petroni et al., 2021) dev 0.05 0.41 0.90 1.70 3.78 6.50 HellaSwag (Zellers et al., 2019) train 3.97 16.60 28.86 46.02 71.68 85.20
train 0.15 0.74 1.42 2.52 5.85 10.49 dev 3.77 15.00 26.31 42.89 69.00 83.51
SQuAD (Rajpurkar et al., 2016) dev 36.94 57.61 64.78 71.31 77.89 81.96 MMLU (Hendrycks et al., 2021b, a) train 3.65 14.14 23.98 38.68 62.22 78.04
test 36.67 57.37 64.50 71.09 77.74 81.84 dev 6.53 21.75 33.64 50.62 75.77 88.18
train 31.84 50.24 57.15 63.33 70.32 74.61 test 6.53 22.01 34.67 51.22 74.35 86.40
QuaRTz  (Tafjord et al., 2019) test 5.87 23.60 38.90 58.67 80.74 90.43 NarrativeQA (s Koˇciský et al., 2018) dev 15.46 23.92 27.10 30.45 35.08 38.89
dev 7.81 26.82 40.62 63.54 84.38 92.19 test 16.80 25.24 28.46 31.50 35.74 38.65
train 7.81 26.82 40.62 63.54 84.38 92.19 train 16.80 25.24 28.46 31.50 35.74 38.65
NQ (Kwiatkowski et al., 2019) dev 23.62 45.60 55.41 64.19 72.51 77.72 OpenBookQA (Mihaylov et al., 2018) dev 3.40 15.20 31.60 50.60 79.00 90.00
test 23.46 45.62 56.32 64.88 74.57 79.72 train 3.35 16.24 28.99 47.85 75.01 88.56
train 22.89 44.97 54.32 62.73 71.75 77.03 test 4.00 16.80 32.20 51.20 74.00 88.80
WebQuestions (Berant et al., 2013) test 19.54 42.86 53.44 63.24 72.34 76.43 Musique (Trivedi et al., 2022) dev 7.36 15.47 21.39 25.98 34.59 41.29
train 21.41 44.71 55.35 64.21 72.71 78.45 train 6.23 13.70 18.54 24.39 33.77 40.81
T-REx (ElSahar et al., 2018; Petroni et al., 2021) dev 45.48 64.34 70.60 76.04 80.42 82.98 TruthfulQA (Lin et al., 2022b) dev 2.33 7.59 12.12 20.69 38.19 56.55
SIQA (Sap et al., 2019) train 0.34 1.52 2.78 4.85 8.92 13.43 HotpotQA (Yang et al., 2018) dev 28.13 45.02 51.92 57.79 65.44 70.67
dev 0.61 2.05 3.58 6.55 11.31 16.58 train 34.89 51.71 58.13 64.09 71.11 75.49
StrategyQA (Geva et al., 2021) train 4.19 15.72 26.29 39.87 52.88 57.86 PopQA (Mallen et al., 2022) test 25.00 38.77 44.70 49.74 56.82 62.36
WOW (Dinan et al., 2019; Petroni et al., 2021) dev 0.20 0.33 0.36 0.39 0.46 0.46 ZSRC (Levy et al., 2017) dev 52.26 70.86 76.29 81.44 85.39 87.73
train 0.26 0.38 0.41 0.45 0.49 0.52 train 50.53 67.23 72.06 76.10 80.35 82.99
Bamboogle (Press et al., 2023) test 5.60 12.80 17.60 24.80 39.20 44.00 EntityQuestions (Sciavolino et al., 2021) test 43.36 60.50 66.08 70.61 75.41 79.06
Table 2. Retrieval performance of BM25 across several selected datasets. The table reports Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Datasets vary in complexity, including open-domain QA (e.g., TriviaQA, NQ), multi-hop reasoning (e.g., 2WikiMultiHopQA, HotpotQA), fact verification (e.g., FEVER, TruthfulQA), and temporal retrieval (e.g., ArchivalQA, ChroniclingQA).
4.2. Retrieval Results
The retrieval performance of BM25 across multiple datasets is shown in Table 2. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Performance varies significantly based on dataset characteristics, with open-domain QA datasets such as TriviaQA and Natural Questions achieving high Top-k accuracy, while more complex multi-hop datasets like 2WikiMultiHopQA and HotpotQA show lower retrieval rates, reflecting the challenges in retrieving supporting evidence across multiple documents. Temporal datasets like ArchivalQA and ChroniclingQA exhibit moderate retrieval performance, likely due to the temporal nature of their content.
For datasets requiring precise fact verification, such as FEVER and TruthfulQA, BM25 struggles, achieving low Top-1 accuracy, indicating the limitations of keyword-based retrieval in distinguishing fine-grained factual nuances. In contrast, reasoning-based datasets like StrategyQA and CommonsenseQA show relatively low retrieval accuracy, highlighting the need for more advanced retrievers capable of capturing implicit reasoning dependencies.
Table 3 presents the retrieval performance of various retrievers across three benchmark datasets: NQ, WebQ, and TriviaQA. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy, providing insights into how different retrieval models rank relevant documents. Among the tested retrievers, dense retrieval methods such as DPR, MSS-DPR, and ColBERT consistently outperform retrievers like MSS and Contriever. MSS-DPR achieves the highest Top-1 accuracy, reaching 50.16% on NQ, 44.24% on WebQ, and 61.63% on TriviaQA, demonstrating the effectiveness of fine-tuned dense retrieval. Standard DPR also performs well, particularly on NQ and TriviaQA, where it reaches 48.67% and 57.47% Top-1 accuracy, respectively. ColBERT, which employs late-interaction ranking, shows competitive performance, particularly in higher recall settings (Top-50 and Top-100).
Retriever Dataset Top-1 Top-5 Top-10 Top-20 Top-50 Top-100
MSS (Sachan et al., 2021) NQ 19.27 41.24 51.27 59.97 69.55 75.56
WebQ 11.66 29.03 39.12 49.21 61.12 68.35
TriviaQA 30.76 52.64 60.52 67.17 74.57 79.11
MSS-DPR (Sachan et al., 2021) NQ 50.16 71.88 77.47 81.44 85.98 88.14
WebQ 44.24 65.00 71.65 76.91 81.84 84.54
TriviaQA 61.63 75.21 79.14 81.85 84.92 86.58
Contriever (Izacard et al., 2021) NQ 22.16 47.28 58.72 67.86 76.01 80.55
WebQ 19.98 43.45 56.39 65.69 74.85 80.11
TriviaQA 34.16 59.48 68.00 73.91 79.83 82.93
DPR (Yates et al., 2021) NQ 48.67 68.78 74.54 79.19 83.71 85.70
WebQ 44.83 65.00 70.62 74.60 78.69 81.64
TriviaQA 57.47 72.40 76.50 79.76 82.96 85.09
ColBert (Santhanam et al., 2021) NQ 42.99 68.78 76.12 82.08 86.15 88.59
WebQ 40.11 64.81 71.56 76.57 81.20 84.50
TriviaQA 57.36 75.90 79.49 82.18 84.74 86.41
Table 3. Comparison of different retrieval models on Natural Questions (NQ), WebQuestions (WebQ), and TriviaQA. The table reports Top-1 to Top-100 accuracy for each retriever.
Reranking/ Model NQ-Test WebQ
Top-1 Top-5 Top-10 Top-20 Top-50 Top-1 Top-5 Top-10 Top-20 Top-50
BM25 - 22.10 43.76 54.45 62.93 72.74 18.80 41.83 52.16 62.40 71.70
UPR (Sachan et al., 2022) T5-small 23.60 49.44 59.97 67.70 75.15 18.55 43.65 55.56 64.91 72.68
T5-base 26.81 52.93 63.32 69.77 76.12 20.66 48.52 58.56 66.48 72.68
T5-large 29.75 55.84 65.67 72.02 76.48 24.21 51.77 60.38 67.27 73.32
T0-3B 35.42 60.44 67.56 72.68 76.75 32.48 56.39 64.17 69.09 73.67
gpt-neo-2.7B 28.75 55.01 64.81 71.13 76.56 24.75 48.91 59.64 66.83 72.63
gpt2 25.95 50.08 60.47 69.05 75.87 20.47 45.32 56.49 65.69 72.78
gpt2-medium 26.75 52.99 63.04 69.80 75.95 22.39 49.01 59.54 66.92 72.44
gpt2-large 26.59 52.35 62.68 70.13 75.95 24.06 49.85 59.84 67.07 72.78
gpt2-xl 27.28 53.85 63.24 70.11 75.84 23.57 50.78 60.48 66.87 72.73
RankGPT (Sun et al., 2023a) llamav3.1-8b 41.55 60.30 66.17 70.60 75.42 38.77 58.36 62.69 67.71 73.12
FlashRank (Damodaran, 2024) TinyBERT-L-2-v2 31.49 53.90 61.57 68.33 74.95 28.54 51.87 60.62 67.42 73.17
MiniLM-L-12-v2 41.02 62.07 67.50 72.49 76.26 37.05 58.07 64.96 70.02 73.96
MultiBERT-L-12 11.99 31.74 43.54 55.01 69.63 12.54 34.30 45.91 55.90 67.91
ce-esci-MiniLM-L12-v2 34.70 57.47 64.81 70.60 76.17 31.84 55.01 62.54 68.60 73.47
RankT5 (Zhuang et al., 2023) base 43.04 62.99 68.47 72.99 76.28 36.95 58.75 64.27 70.07 74.45
large 45.54 64.98 70.02 73.85 76.81 38.77 60.28 66.48 70.96 74.31
3b 47.17 66.45 70.85 74.18 76.89 40.40 60.33 66.58 71.16 74.45
Inranker (Laitz et al., 2024) small 15.90 40.11 46.84 56.23 69.83 14.46 36.36 46.25 57.48 69.98
base 15.90 40.69 48.11 56.62 69.66 14.46 37.45 46.80 56.54 69.68
3b 15.90 40.22 48.06 57.97 69.00 14.46 37.30 46.11 57.48 69.34
LLM2Vec (BehnamGhader et al., 2024) Meta-Llama-31-8B 24.32 49.22 59.55 68.08 75.26 26.72 50.39 60.48 68.40 73.47
MonoBert (Nogueira et al., 2019) large 39.05 61.71 67.89 72.63 76.56 34.99 57.33 64.56 70.12 73.96
Twolar (Baldelli et al., 2024) twolar-xl 46.84 65.29 70.22 73.62 76.86 41.68 60.82 67.07 71.55 74.40
Echorank (Rashid et al., 2024) flan-t5-large 36.73 55.09 59.11 62.38 62.38 31.74 53.69 58.75 61.51 61.51
flan-t5-xl 41.68 55.15 59.05 62.38 62.38 36.22 52.55 57.18 61.51 61.51
Incontext Reranker (Chen et al., 2024) llamav3.1-8b 15.15 38.50 57.11 69.66 76.48 18.89 41.83 52.16 62.40 71.70
Lit5 (Tamber et al., 2023) LiT5-Distill-base 40.05 59.80 65.95 70.80 75.73 36.76 56.79 63.48 67.81 73.12
LiT5-Distill-large 44.40 62.57 67.59 71.74 76.01 39.66 58.02 64.56 68.94 73.67
LiT5-Distill-xl 47.81 64.07 68.55 72.07 76.26 42.37 59.44 65.55 69.38 73.62
LiT5-Distill-base-v2 42.57 61.49 66.73 70.83 75.56 39.61 57.97 64.22 68.55 73.32
LiT5-Distill-large-v2 46.53 63.10 67.83 71.91 75.87 41.97 59.49 65.64 68.65 72.98
LiT5-Distill-xl-v2 47.92 64.18 69.03 72.40 76.17 41.53 60.13 65.69 69.04 73.27
Sentence Transformer Reranker all-MiniLM-L6-v2 (Wang et al., 2020) 33.35 57.34 65.37 70.72 76.01 30.95 53.44 62.10 68.01 73.52
GTR-base (Ni et al., 2021) 39.41 59.08 65.95 71.32 76.03 36.56 58.21 64.32 69.14 73.62
GTR-large 40.63 62.07 68.25 72.32 76.73 38.97 58.51 65.30 70.02 73.57
GTR-xl 41.55 61.82 67.78 72.65 76.81 38.92 59.69 66.04 70.57 74.01
GTR-xxl 42.93 63.32 68.55 73.01 77.00 39.41 60.03 65.89 70.47 74.01
T5-base (Raffel et al., 2020) 31.19 55.84 63.60 70.22 76.06 29.77 54.92 62.84 68.70 73.52
T5-xxl 38.89 61.46 67.78 72.49 76.64 35.82 58.26 65.206 70.32 74.01
T5-large 30.80 55.59 63.35 70.02 76.37 30.51 55.06 61.71 68.35 73.37
Table 4. Performance of re-ranking methods on BM25-retrieved documents for Natural Questions (NQ-Test) and WebQuestions (WebQ-Test). Results are reported in terms of Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy, highlighting the impact of various re-ranking models on retrieval effectiveness.
4.3. Re-Ranking Results
The performance of various re-ranking methods applied to BM25-retrieved documents is presented in Table 4. The table reports Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy for two datasets: Natural Questions (NQ-Test) and WebQuestions (WebQ). The baseline BM25 retrieval scores serve as a reference, demonstrating the extent to which re-ranking improves document ranking quality across different ranking models.
Across all models, re-ranking consistently improves retrieval effectiveness, with notable gains observed in Top-1 accuracy. Transformer-based ranking models such as UPR (T5-based models), RankGPT, FlashRank (MiniLM and TinyBERT), and RankT5 achieve substantial improvements over the BM25 baseline. In particular, RankT5-3B outperforms other ranking models, achieving a Top-1 accuracy of 47.17% on NQ-Test and 40.40% on WebQ, highlighting the effectiveness of T5-based ranking models for large-scale retrieval. RankGPT, leveraging LLaMA-3.1-8B, also demonstrates strong performance, particularly in WebQ, where it improves Top-1 accuracy from 18.80% (BM25) to 38.77%.
Among lighter-weight models, FlashRank (MiniLM and TinyBERT) and MonoBERT-large provide solid improvements while maintaining efficiency. MiniLM-L-12-v2 achieves 41.02% Top-1 accuracy on NQ-Test and 37.05% on WebQ, making it a strong candidate for deployment scenarios where computational efficiency is critical. The MonoBERT-large model reaches 39.05% on NQ-Test and 34.99% on WebQ, reinforcing the effectiveness of BERT-based cross-encoders for document ranking.
Models that integrate listwise and contrastive ranking techniques, such as Twolar, and LiT5 re-rankers, show competitive performance. Twolar-xl achieves 46.84% Top-1 accuracy on NQ-Test and 41.68% on WebQ, while LiT5-Distill-xl reaches 47.81% on NQ-Test and 42.37% on WebQ, demonstrating that fine-tuned T5 models can significantly improve ranking quality. The sentence-transformer rerankers, particularly GTR-xxl, achieve 42.93% and 39.41% Top-1 accuracy on NQ-Test and WebQ, respectively, making them effective in ranking diverse retrieval outputs.
Overall, the results highlight the importance of re-ranking in improving retrieval effectiveness. While BM25 provides a strong initial retrieval, the application of neural re-rankers significantly enhances retrieval quality, particularly for Top-1 and Top-5 results. The effectiveness of different models varies based on dataset characteristics, with larger transformer-based models excelling in ranking accuracy, while lighter-weight models offer a balance between accuracy and efficiency.
5. Conclusion
We introduced Rankify, a modular toolkit that unifies retrieval, re-ranking, and retrieval-augmented generation (RAG) within a cohesive framework. By integrating diverse retrievers, state-of-the-art re-rankers, and seamless RAG pipelines, Rankify streamlines experimentation and benchmarking in information retrieval. Our experiments demonstrate its effectiveness in enhancing ranking quality and retrieval-based text generation.
Designed for extensibility, Rankify allows for easy integration of new models and evaluation methods. Future work will focus on improving retrieval efficiency, optimizing re-ranking strategies, and advancing RAG capabilities. The framework is open-source, actively maintained, and freely available for the research community.